Setrakian Industries needs to raise $97.1 million to fund a new project. The company will sell bonds that have a coupon rate of 6.27 percent paid semiannually and that mature in 20 years. The bonds will be sold at an initial YTM of 6.79 percent and have a par value of $2,000. How many bonds must be sold to raise the necessary funds? (Round your intermediate calculations to two decimal places and final answer to the nearest whole number.)
Answer #1

The value of the bond is calculated using the PV function as follows:

=PV(rate,nper,pmt,fv)

=PV(6.79%/2,20*2,6.27%/2*2000,2000)

=1887.12

No of bonds to raise the necessary cash=97,100,000/1887.12 = 51454 bonds